two cars start at the same moment from opposite stations 7 miles apart. they drive toward each other on the same road at constant speeds, one at 25mph and the other at 45mph. after how many minutes will they pass each other on the road
Answer by mananth(16946)

Car I speed = 25
Car II speed = 45
they are driving towards each other
add up speeds
effective speed = 70 mph
distance = 7 miles
t= 7/70
t=1/10 hour
1/10 * 60 = 6 minutes